### v1

# Document Type Classification using File Names

Abstract
Rapid document classification is critical in several time-sensitive applications like digital forensics and large-scale media classification. Traditional approaches that rely on heavy-duty deep learning models fall short due to high inference times over vast input datasets and computational resources associated with analyzing whole documents. In this paper, we present a method using lightweight supervised learning models, combined with a TF-IDF feature extraction-based tokenization method, to accurately and efficiently classify documents based solely on file names that substantially reduces inference time. This approach can distinguish ambiguous file names from the indicative file names through confidence scores and through using a negative class representing ambiguous file names. Our results indicate that file name classifiers can process more than 80% of the in-scope data with 96.7% accuracy when tested on a dataset with a large portion of out-of-scope data with respect to the training dataset while being 442.43x faster than more complex models such as DiT. Our method offers a crucial solution for efficiently processing vast datasets in critical scenarios, enabling fast, more reliable document classification.
Document Type Classification using File Names
Zhijian Li     Stefan Larson     Kevin Leach Vanderbilt University {firstname.lastname}@vanderbilt.edu
1 Introduction
Classifying massive amounts of documents at scale is often an important yet challenging task, especially at industrial scale, where applications of document classification include sensitive data identification, document review for legal cases, spam and malware flagging, etc. For example, in sensitive data identification, document sensitivity risk scores can be assigned according to type, e.g., academic_transcript $\rightarrow$ sensitive, medical_intake_form $\rightarrow$ sensitive, press_release $\rightarrow$ low-risk, etc.
Typical approaches to document type classification use text-, image-, layout-, and multimodal classifiers (e.g., Li et al. (2022); Huang et al. (2022); Pham et al. (2022)), but these models can be prohibitively slow at the scale required in many industry settings, as these models must execute sophisticated operations on the contents of a very large quantity of documents. Thus, there is a need for techniques that speed up the document classification pipeline to enable scale.
In this paper, we introduce and investigate the task of classifying documents by type (e.g., resume, press_release, etc.) based on their file names. Figure 1 displays several examples where documents can be categorized based on their file names. An added challenge is to correctly identify cases that are ambiguous. Since many file names are indicative of their contents (e.g., john_resume_2024.pdf $\rightarrow$ resume), we hypothesize that much simpler classifiers can be trained and used to categorize these documents.
Figure 1: Overview of the file name classification task. An ideal classifier can categorize indicative file names correctly, while also being able to defer predictions on ambiguous file names.
Our experimental result shows that the Random Forest file name classifier with a Trie tokenizer can achieve 99.6% accuracy on the indicative file names and 96.7% accuracy on a dataset distinct from the training dataset with a large portion of out-of-scope documents while deferring less than 80% of the in-scope documents. This Random Forest file name classifier requires only $1.23\times 10^{-4}$ seconds per prediction which is 442.43x faster than DiT Li et al. (2022), an exemplar image-based deep learning model that requires access to each document.
File Name Indicative or Ambiguous Type
Cerno-Woodshop-Assistant-Opportunity.pdf indicative job_posting
TcallsenResume2020v0710-aem-react.pdf indicative resume
The-Willow-Restaurant-Menu-0422.pdf indicative restaurant_menu
StJudeNewsRelease2011.pdf indicative press_release
MX_7090N_8090N_Spec_Sheet.pdf indicative specification
dear-colleague-letter-fda-samhsa.pdf indicative letter
BACB_September2023_Newsletter-231201-a.pdf indicative newsletter
Intermolecular-USCAN.pdf ambiguous press_release
MO_Farmington.pdf ambiguous restaurant_menu
12042023-NMPH.pdf ambiguous job_posting
plnbudm3-30-4-6s12.pdf ambiguous meeting_minutes
5bba5a94c8f1454f98d7b01be13289b9.pdf ambiguous policy
Table 1: Examples of indicative file names and ambiguous file names.
2 Related Work
In this section we discuss prior work on classifying files using file names, and document type classification.
2.1 Classification with File Names
Document type classification via file names is an understudied topic. While prior research on categorizing content via file names has shown promise, it has largely focused on binary classification. For instance, Al-Nabki et al. (2023) proposed an approach to classify Child Sexual Exploitation Material (CSEM) through analyzing the file names and file paths of files when law enforcement officers seize a computer from a CSEM suspect due to the time-consuming nature of traditional and heavy weight models. Similarly, Jáñez Martino et al. (2021) accurately detected spam emails through analyzing email names using a Naive Bayes classifier, and Nguyen et al. (2019) demonstrated that a character-level CNN can accurately classify file names that contain enough information indicative of malware. Additionally, Sánchez-Paniagua et al. (2022) accurately detected phishing URLs using a logistic regression model combined with Token Frequency-Inverse Document Frequency (TF-IDF) feature extraction.
These studies demonstrate that binary classification through file names alone can be effective without deeper analysis of the contents of the files themselves. However, multi-class file name classification tasks, such as document type classification, is still under-researched. Moreover, the issue of handling file names that fall outside the the scope of the classifiers or those that lack clear indicators for accurate classification has not yet been adequately addressed in the existing literature.
2.2 Document Type Classification
Prior work on document type classification includes the development of large image-, text-, and page layout-based models for categorizing documents, mostly from datasets derived from the IIT-CDIP tobacco corpus (e.g., RVL-CDIP Harley et al. (2015) and Tobacco3482 Kumar and Doermann (2013)). These large models include the LayoutLM family (LayoutLM Xu et al. (2020), LayoutLMv2 Xu et al. (2021), LayoutLMv3 Huang et al. (2022)), DiT Li et al. (2022), DocFormer Appalaraju et al. (2021), Donut Kim et al. (2022), and LayoutLLM Fujitake (2024), etc. However, these models are costly to train and costly to run. Image-based models like CNNs and Transformers must operate on images of document pages, and text- and layout-based models must run on variants of the document from which text has been extracted via Optical Character Recognition (OCR), which adds a layer of complexity for each document that must be classified.
3 Problem Overview
In this study, we examine whether we can efficiently and accurately classify documents only based on their file names. Table 1 shows examples of file names and their categories. To enhance the clarity of our analysis, we distinguish the file names into the two following broad categories:
  • Indicative: File names that contains the category name, a synonym of that category, or an abbreviation of the category name. Examples of indicative file names include john_smith_resume_2024.pdf (a resume document), john_smith_cv_2024.pdf (a resume document, where “cv” is used as a synonym for “resume”), and RTKRequestForm.pdf (a form document).
  • Ambiguous: File names that does not contain the category name itself or a synonym of the category name. Examples of ambiguous file names include 20-005-10.pdf and document.pdf. We note that ambiguous file names are—to borrow terminology from Larson et al. (2019)—“out-of-scope” with respect to the in-scope or in-domain file type categories defined in the dataset.
A key aspect of our approach is to determine whether the file name is indicative based on the confidence score associated with the predicted class for that file name. In cases where the confidence is too low, we defer the file name. In practice, we would instead use a more complex model at increased computational cost if the file is deferred by the filename classifier. In cases where the confidence score is high enough, we desire a highly accurate class prediction to forego the additional expense of running a deeper, more complex model involving the file’s contents. With a higher confidence threshold, the file name classifiers are more accurate with these predictions, but predicts a smaller portion of all the file names, ultimately adding computational burden to the heavier-duty model in the next stage. We discuss the trade-off and optimal confidence threshold for each file name classifier in Section 6. We also measure the AUROC of the confidence scores between indicative and ambiguous file names to analyze if the file name classifiers can produce confidence scores that differentiate between indicative and ambiguous file names.
In this study, we only evaluate the inference speed of the heavier-duty models as we are interested how much faster the file name classifiers are compared to models such as DiT, rather than trying to outperform the more complex models in accuracy. This study aims to address whether file name classifiers can accurately predict indicative file names. Additionally, we evaluate if file name classifiers predict a large portion of the in-scope data with high accuracy when tested on a dataset containing indicative, ambiguous, and out-of-scope file names.
As shown in Li et al. (2024), adding a class of out-of-scope data in training improves classifier robustness when identifying out-of-scope data from the in-scope data. We hypothesize that by populating a new category with the ambiguous file names, coupled with marking predictions for supported categories with low confidence scores as ambiguous file names, the file name classifiers’ prediction accuracy can be improved.
4 Method
In this section, we discuss our approach to leveraging indicative and ambiguous file names to develop a fast, accurate document classification scheme that reduces the overall computation resources required to classify large volumes of documents. Our approach consists of a tokenization scheme for splitting input file names as well as training and evaluating light-weight machine learning pipeline.
4.1 Tokenization
Figure 2: An example of the Universal Tokenizer tokenizes a fileame with the Trie Tokenizer and BERT Autotokenizer.
We apply a Universal tokenizer to clean all input file names. Because many file names are gathered from URLs, we first each file name by replacing URL escape sequences (e.g., “%xx”) with the relevant single-character equivalent. We then split the file names by all punctuation characters, white space, and camel case boundaries (i.e., when a lower-case character is immediately followed by an upper-case character).
The Universal Tokenizer is not equipped to handle cases where file names contain words joined together without any punctuation or camel case such as "coursecatalog.pdf".
From randomly sampling 1,000 file names and annotating the Universal tokenizer’s performance, we discover that 3.4% file names needs further tokenization.
Figure 2 shows an example where the Universal tokenizer fails to separate “Coursesyllabus” into “Course” and “syllabus,” and additional processing is needed to correctly tokenize this file name.
To improve the tokenization on those file names, we first calculate the Token Frequency-Inverse Document Frequency of every token in relation to the category. After lemmatizing the output tokens with NLTK’s WordNetLemmatizer Bird and Loper (2004), tokens with TF-IDF higher than a threshold $k$ are identified as “keywords.” We choose $k$ to be 0.1 in this study. The trade-offs of selecting different value of $k$ is discussed in Section 6.5. The keywords are then added to a trie data structure to create a Trie tokenizer that can effectively locate and split important words that are joined together if the words in the file names are not separated with special characters, whitespace, or camel case. Another approach involves adding the "keywords" to a BERT Autotokenizer Wolf et al. (2020) using the “add_tokens” function in the Hugging Face library. For example, the Trie or BERT tokenizer will split file names such as “coursecatalog” in “course” and “catalog” to enable effective bag-of-words encoding.
This leaves us with a set of tokens for each file name, allowing us to find indicative cases where the file name makes a clear and straightforward hint as to the underlying document’s class.
4.2 Classification
We use the following text classifiers in our experiments:
Support Vector Machine (SVM), a supervised learning model that classifies data by finding a separating hyperplane that maximizes the margin. We use the linear kernel and the regularization parameter c = 0.1 as this combination yields the highest accuracy during our hyperparameter search where we compare the accuracy on the SVMs trained and tested on indicative file names with different kernels and value of c. The SVM model is implemented with Scikit-Learn which uses a one-vs-all approach.
Naive Bayes, a supervised learning model that learns the distribution of inputs by assuming that all the features are independent of each other. We implement this model with Scikit-Learn.
Random Forest, a supervised learning model that aggregates the prediction of many decision trees to prevent over-fitting. We implement this model with 100 decision trees through Scikit-Learn Pedregosa et al. (2011).
eXtreme Gradient Boosting (XGBoost), a supervised learning model that is implemented with gradient boosted decision trees. It is implemented with the XGBoost library Chen and Guestrin (2016).
BERT, a neural network based on a transformer model, capture the context of each word Devlin et al. (2019). In this study, we fine-tune the BERT-base-uncased model from Hugging Face [^1] https://huggingface.co/google-BERT/BERT-base-uncased on the Web Search dataset to classify file names.
When preprocessing the file names, we first use the tokenization technique discussed in Section 4.1. Subsequently, for the SVM, Naive Bayes, Random Forest, and XGBoost classifiers, we encode the file name tokens with a bag-of-words encoding. For BERT, we join the tokens with space and then encode the string with the BERT tokenizer.
4.3 Metrics
For a selected confidence threshold $t$, we analyze the prediction accuracy by calculating #$\frac{\#\mathrm{correct}}{\#\mathrm{correct}+\#\mathrm{incorrect}}$, the number of correct predictions divided by the number of file names whose prediction confidence exceeds $t$. We examine the prediction rate by calculating #$\frac{\#\mathrm{correct}+\#\mathrm{incorrect}}{\mathrm{total}\#\mathrm{documents}}$, the sum of the Correct and Incorrect predictions over the total number of file names.
To analyze the models’ ability to distinguish ambiguous file names from the indicative file names, we calculate the confidence score for each prediction. For BERT, we use the highest softmax confidence score for each prediction Hendrycks and Gimpel (2016). For SVM, Naive Bayes, and Random Forest, we use the highest value from the “predict_proba” function through Scikit-Learn. Through calculating the Area Under the Receiver Operating Characteristic curve (AUROC) between the indicative file names and the ambiguous file names, we can measure how well the models can distinguish the ambiguous file names through confidence scores and avoid mispredicting file names that do not contain enough information. We note that AUROC has been used in prior work in the document classification space (e.g., Larson et al. (2022)).
5 Datasets
In this section we introduce two datasets used as part of our evaluation.
5.1 Web Search Dataset
We collected a dataset with 1,235 PDF files from 15 categories through web searches. A list of these 15 categories is shown in Table 6 in the Appendix.
We considered search prompts of the format: <category_name> filetype:pdf (where filetype:pdf specifies that we desire PDF documents from the web search engine). For example, we gather the PDF files for restaurant menus by searching restaurant menu filetype:pdf. Then, we manually open the files to verify that they belong to the desired category, and then record the verified document’s file name. We also note that synonyms or related words and phrases are also used when searching for category names. For instance, when searching for restaurant menus, we also searched for “diner menu,” “bar menu,” etc.
The Web Search dataset contains 15 categories, which is similar in number to the RVL-CDIP dataset Harley et al. (2015) which contains 16 categories. Table 1 displays more examples of indicative and ambiguous file names from this dataset. After manual annotation, we found that our Web Search dataset contains 978 indicative file names and 257 ambiguous file names. This breakdown of file names is used in our methodology to establish a “best-case” performance — for indicative file names, we desire a shallow, low-cost model that accurately classifies the given document into the best predicted category. The ambiguous file names that remain do not contain enough information to successfully classify, thus in practice we want the file name classifiers to defer to the heavier-cost model to analyze the content of those documents.
5.2 Common Crawl Dataset
In our evaluation, we also consider another open source dataset to demonstrate generalizability of our approach. We use a subset of the CC-MAIN-2021-31 dataset, which consists of PDFs, URLs, and metadata of eight million documents collected by Common Crawl from the web in July and August of 2021.222https://digitalcorpora.org/corpora/file-corpora/cc-main-2021-31-pdf-untruncated/ We randomly selected 1,198 documents that contain valid URLs and are in English. Of these 1,198 file names, there are 266 file names belonging to one of the 15 document type categories defined in the Web Search dataset from Section 5.1, and the remaining 932 file names belong to out-of-scope or categories unknown to the classifier. This distribution of categories for the file names differs from that of the Web Search dataset, and can simulate the robustness of file name classifiers trained on datasets that contain fewer categories and are smaller than the datasets they are tasked to predict when not every category is of interest to be classified.
The correct category of the Common Crawl dataset are obtained by manual annotation by two annotators without considering whether the file name is indicative or ambiguous. When the annotators cannot reach an agreement on the appropriate type, the document is discarded.
Models Without Negative Class With Negative Class
Tokenizer Threshold Accuracy Rate Tokenizer Threshold Accuracy Rate
SVM Trie 0.18 0.933 0.837 BERT 0.50 0.994 0.807
Naive Bayes BERT 0.15 0.893 0.877 BERT 0.22 0.979 0.803
Random Forest Trie 0.24 0.941 0.829 Trie 0.62 0.996 0.805
XGBoost Trie 0.13 0.898 0.895 Trie 0.60 0.993 0.809
BERT Trie 0.32 0.948 0.803 Trie 0.82 0.992 0.835
Table 2: The confidence threshold value and tokenizer combination that maximizes the prediction accuracy for every model trained with or without a negative class while the prediction rate is above 0.8 for all confidence threshold between 0 and 1 incremented by 0.01, trained and tested on the Web Search dataset.
6 Experiments
We discuss the experimental results and settings in this section.
SVM, Naives Bayes, Random Forest, XGBoost, and BERT classifiers are trained on only the indicative file names from the Web Search dataset, and evaluated against the rest of the data from the Web Search dataset and the Common Crawl dataset. In an additional experimental set up, all the classifiers are trained on indicative file names and ambiguous file names (the latter being a single additional classification class) from the Web Search dataset and evaluated against both datasets. We perform 5-fold cross validation and calculate the average prediction accuracy and prediction rate of different setups (defined in Section 4.3). The subsections below discuss the accuracy and speed of our file name classifiers, as well as their ability to distinguish ambiguous data.
6.1 Model Accuracy for Indicative File Names
First, we show that our approach can accurately classify indicative file names. Table 2 shows the confidence thresholds and tokenizer combination that maximizes the prediction accuracy of every model while the prediction rate is above 0.8 when tested on only indicative data. Predictions that result in a confidence score lower than the threshold or produce an ‘ambiguous’ label are deferred by the file name classifier. With a prediction rate of 0.805, the Random Forest file name classifier with a Trie tokenizer trained with a negative class produces the highest prediction accuracy of 99.6% showing that this file name classifier can produce accurate predictions for the majority of the indicative data while deferring the rest. SVM, Random Forest, XGBoost, and BERT when trained with a negative class, all achieved above 99% accuracy with a prediction rate above 0.8. When the negative class is not used in training, every model’s prediction accuracy decreases substantially, with the two most accurate classifiers with BERT and Random Forest, achieved 94.8% and 94.1% accuracy, respectively. We observed insubstantial differences in prediction accuracy for every file name classifier with different tokenizers.
Figure 3: The confidence score distribution for Random Forest classifier with Trie tokenizer on indicative and ambiguous file names from the Web Search dataset.
Tokenizer SVM Naive Bayes Random Forest XGBoost BERT
Trie 0.907 0.887 0.930 0.894 0.882
BERT 0.916 0.891 0.927 0.896 0.878
Table 3: The AUROC of confidence scores between the indicative and ambiguous file names for all the classifiers with a Trie or BERT tokenizer.
Models Without Negative Class With Negative Class
Tokenizer Threshold Accuracy Tokenizer Threshold Accuracy
SVM Trie 0.60 0.748 BERT 0.46 0.931
Naive Bayes BERT 0.59 0.729 BERT 0.29 0.782
Random Forest BERT 0.61 0.934 BERT 0.65 0.967
XGBoost Trie 0.83 0.879 BERT 0.71 0.915
BERT Trie 0.97 0.866 BERT 0.97 0.842
Table 4: The confidence threshold value and tokenizer that maximizes the prediction accuracy for every model trained with or without a negative class while the prediction rate is above 0.178 for all confidence threshold between 0 and 1 incremented by 0.01. The models are trained on the Web Search dataset and evaluated on the Common Crawl dataset.
6.2 Detection Ability for Ambiguous File Names
We evaluate our file name classifiers against the ambiguous file names and evaluate our model’s robustness predicting ambiguous file names by calculating the AUROC of the confidence scores between the indicative and ambiguous file names. Figure 3 shows minor overlap between the confidence scores for the indicative and ambiguous file names for the Random Forest file name classifier with the Trie tokenizer.
Table 3 displays the AUROC of the confidence scores between indicative and ambiguous file names for all models and each tokenizer. The Random Forest file name classifier with the Trie tokenizer achieves the highest AUROC which shows that this classifier produces confidence scores that best differentiate indicative and ambiguous file names. The differences in AUROC for different tokenizers are also not substantial for all models.
Due to the differences in how we obtain confidence scores for each models, the optimal confidence threshold differs between models. Increasing the confidence threshold increases the prediction accuracy while decreasing the prediction rate. In practice, we must select an appropriate confidence threshold to balance the tradeoff between the prediction accuracy and the fraction of documents accurately predicted as indicative — the documents that are deferred as ambiguous would require further examination using a deeper and more complex model. Thus, the average time and resource expenditure per document classified would increase with the increase in confidence threshold for predicting a document file name as indicative.
Figure 4 shows the prediction accuracy and prediction rate of the Random Forest classifier on all the file names in the Web Search dataset (both indicative and ambiguous) for for all confidence thresholds between 0 and 1, in intervals of 0.01 per iteration. Figure 5 shows the overall accuracy and overall time per document required for a Random Forest classifier with the Trie tokenizer trained with a negative class and a DiT model to process all the deferred predictions from the file name classifier tested on the Web Search data including indicative and ambiguous file names. With the assumption that all DiT predictions are correct, overall accuracy is calculated by prediction accuracy $\times$ prediction rate $+$ ($1-$ prediction rate), and overall time is calculated by average inference time of Random Forest + average inference time of DiT * ($1-$ prediction rate). The inference time for each model is discussed in Section 6.4. At a confidence threshold of 0.66, the improvement in prediction accuracy begins to taper off, while the decrease in prediction rate accelerates, and the overall time increases more steeply. This indicates that 0.66 is an optimal confidence threshold for the Random Forest file name classifier.
Figure 4: Prediction Accuracy vs Prediction Rate at all confidence score threshold between 0 and 1 in increments of 0.01 for a Random Forest classifier with the Trie tokenizer trained with a negative class predicted on all the file names in the Web Search dataset including ambiguous file names.
Figure 5: Overall Accuracy vs Overall Time (seconds) at all confidence score threshold between 0 and 1, incremented by 0.01 for a Random Forest classifier with the Trie tokenizer trained with a negative class predicted on all the file names in the Web Search dataset including ambiguous file names, and DiT model to process the documents deferred by the Random Forest model.
6.3 Evaluation on Common Crawl Dataset
We train file name classifiers on the Web Search dataset described in Section 5.1 and evaluate our models against the Common Crawl dataset described in Section 5.2.
Recall that 22.2% of the file names in our Common Crawl dataset belong to a category in the Web Search dataset. Without further annotating whether each of those in-scope file names are indicative, we evaluate how accurate our file name classifiers are while predicting at least 80% of the in-scope data, which is 17.8% of all the documents. Table 4 displays the confidence threshold value and the tokenizer that maximizes the prediction accuracy while achieving a prediction rate of at least 0.178. The result displays that the Random Forest classifier with a BERT tokenizer achieves 96.7% prediction accuracy when trained with a negative class, and is the most robust classifier at predicting the Common Crawl dataset. The SVM file name classifier trained with a negative class and the Random Forest classifier trained without a negative class also reached above 93% prediction accuracy with the BERT tokenizer.
The results indicate that the Random Forest file name classifier trained with a negative class can accurately predict the majority of the in-scope documents from a distinctive dataset with a large portion of out-of-scope documents.
SVM Naive Bayes Random Forest XGBoost BERT
-10.34 -13.34 -8.95 -10.82 -2.79
Table 5: Logarithm of average inference time for SVM, Naives Bayes, Random Forest, XGBoost, and BERT file name classifiers, calculated by applying the natural log to the average inference time in seconds.
6.4 Inference Time
Table 5 shows the file name classifiers’ average log inference time per prediction for each model on a computer with a single 24-core x84_64 CPU and no GPUs.
Among all the models, Naive Bayes was the fastest, with a average log inference time of $-13.34$, requiring $1.61\times 10^{-6}$ seconds per prediction. The BERT file name classifier’s inference time is substantially slower than all four other models, requiring $6.14\times 10^{-2}$ seconds per prediction. The Random Forest classifier that achieved the highest prediction accuracy on both the Web Search dataset and the Common Crawl dataset requires $1.30\times 10^{-4}$ seconds per prediction.
In comparison, when used on the same computer as the file name classifiers, the DiT model on average requires $0.0574$ seconds per image, and the BERT model that uses the first 512 tokens of a document as input requires $0.0678$ seconds per document, both by analyzing the content of the documents instead of the file names. The robust Random Forest file name classifier is 442.43x faster than the DiT model even if the DiT model only processes the first page of every document and is 552.60x faster than the BERT model that processes the content of the documents, offering a dramatic speedup.
6.5 Trade-offs for $k$ Value Selection
Figure 6: Prediction Accuracy vs Prediction Rate at 0.8 confidence threshold for the Random Forest file name classifier with a BERT tokenizer using various k-values on indicative file names from the Web Search dataset.
In Section 4.1, we mentioned that tokens with TF-IDF larger than $k$ are considered "keywords" and added to the tokenizer. We examine the trade-offs for the selection of $k$ value for the BERT tokenizer by comparing the prediction accuracy and prediction rate with various $k$ values. The previous subsections indicate that the Random Forest file name classifier produces the highest prediction accuracy with a fixed minimum prediction rate. Figure 6 displays the prediction accuracy and prediction rate at 0.8 confidence threshold for the Random Forest file name classifiers with $k$ values from 0 to 0.5. For $k$ values between 0.04 and 0.43, the prediction rate remains steady around 0.865 while the prediction accuracy is the highest for $k$ values between 0.025 to 0.132. Thus, a $k$ value should be selected between 0.025 and 0.132.
7 Conclusion
We present a new approach to multi-class document classification that dramatically improves classification efficiency while achieving high accuracy. Our evaluation shows that a Random Forest classifier, when tasked to predict at least 80% of the in-scope data, can achieve 99.6% prediction accuracy on indicative file names when trained and tested on datasets with the same category distribution with a Trie tokenizer. It also achieve 96.7 % prediction on the in-scope file names when trained and tested on data with different distribution of categories with a BERT tokenizer. The Random Forest classifier with a Trie tokenizer can also produce confidence scores that can effectively distinguish indicative and ambiguous file names, resulting in an AUROC of 0.93 for confidence scores between indicative and ambiguous. Meanwhile, the Random Forest classifier is 442.43x faster than the DiT model, indicating that uses the Random Forest file name classifier can provide a substantial overall speedup in document classification. We hope that our technique and analysis will enable faster and computationally efficient document classifiers without sacrificing on accuracy.
Future Work
In this study, we solely focused on document classification using file names and obtained promising results. We believe file metadata — such as file size, page count, number of images, etc. — can be used along with file names to improve the classifier accuracy without substantial increase in inference time. Future work can investigate the potential benefits of combining metadata with file name to improve classification accuracy and or handle ambiguous data or out-of-scope data more effectively.
Appendix A Appendix
Table 6 lists the categories used in the Web Search dataset.
Dataset Availabilty.
We provide both the Web Search and Common Crawl datasets in an anonymous Github repository: https://github.com/frank7li/Document-Type-Classification-using-File-Names.
Category
specification
resume
restaurant_menu
price_list
press_release
policy
newsletter
meeting_minutes
map
letter
job_posting
guide
form
course_syllabus
bill_act_law
Table 6: Document type categories used in the Web Search dataset.